OpenStudy (naga447):

Medal for who ever helps me!!! (graphing) What would the solution to these two equations be? y - 2x = -1 x + 3y = 4

OpenStudy (anonymous):

how fast you want it ?

OpenStudy (anonymous):

i mean "explanation" or how i would get the answer in three seconds?

OpenStudy (naga447):

As fast as you can explain it to me lol

OpenStudy (naga447):

I'd like to know what I'm doing

OpenStudy (anonymous):

\[y-2x=-1\\ x+3y=4\] rewrite so they line up \[-2x+y=-1\\ x+3y=4\]

OpenStudy (anonymous):

now the x over the x and the y over the y multiply the second equation by 2 \[-2x+y=-1\\2x+6y=8\]

OpenStudy (anonymous):

add in a column so the x terms go bye bye leaving \[7y=7\] which makes \(y=1\)

OpenStudy (naga447):

Okay that helped a lot! having someone go though the steps! Thank you very much~

OpenStudy (anonymous):

yw once you have \(y=1\) i hope it is clear how to find \(x\)

OpenStudy (naga447):

I I just do basically the same thing except I'd be solving for x

OpenStudy (anonymous):

no no you don't need all that

OpenStudy (anonymous):

since \[x+3y=4\] and \(y=1\) you have \[x+3=4\] making \(x=1\)

OpenStudy (naga447):

Oh lol that's much simpler than redoing all of that thank you again!

OpenStudy (anonymous):

sure is ! yw